# Behavior-Inspired Neural Networks for Relational Inference

### v2

@@ 4.2 Latent nonlinear opinion dynamics
We use the nonlinear opinion dynamics formulation in Equation (3.1) to model the evolution of agent preferences on a set of latent categories,
@@
We learn the intrinsic agent parameters, $\mathbf{d}$, $\mathbf{u}$, and $\boldsymbol{\alpha}$, the extrinsic parameter $\mathbf{A}^{\mathrm{o}}$, and compute the inter-agent communication matrix $\mathbf{A}^{\mathrm{a}}_{t}$ at each timestep $t$.
@@
where $\Delta t$ is dataset dependent.
@@ 4.3 Communication matrix
We define the communication matrix $\mathbf{A}^{a}_{t}$ as a function of inter-agent proximity. In our baseline BINN model, we define $\mathbf{A}^{a}_{t}$ as the distance between observed positions,
$a^{a}_{(i,j),t}=\lVert\mathbf{p}_{i,t}-\mathbf{p}_{i,t}\rVert^{2}.$ (10)
We also introduce Behavior-Inspired Neural Network+ (BINN+), where we incorporate an informed hypothesis of how agents in the system communicate. For example, for systems in which agents have more influence on each other when they are closer (e.g., human interaction) we define $\mathbf{A}^{a}_{t}$ as the inverse distance between observed positions,
$a^{a}_{(i,j),t}=\frac{1}{\lVert\mathbf{p}_{i,t}-\mathbf{p}_{i,t}\rVert^{2}+\epsilon}$ (11)
where $\epsilon$ is small.
@@ 4.4 Decoder
Our decoding network $D_{x}$, is an MPNN [54] that maps agent preferences to predictions of agent states. At every timestep $t$, the latent preferences of the multi-agent system are represented as a fully-connected graph with node values determined by the preferences of each agent.
Formally, our decoder $D_{x}$ performs the following message passing functions for agent $i$ at timestep $t$:
@@ 4.5 Loss function
$\mathcal{L}_{\mathrm{latent}}=\frac{1}{T\mathcal{N}_{a}}\sum_{t=1}^{T}\sum_{i=1}^{\mathcal{N}_{a}}\Big{\lVert}\mathbf{z}_{i,t+1}-f_{\mathrm{NOD}}\left(\mathbf{z}_{i,t}\right)\Big{\rVert}^{2},$ (18)
and encourages alignment between the representations encoded by $E_{z}$ and those predicted by $f_{\mathrm{NOD}}$.
@@ 5.1 Interpretability of the latent space
Figure 4: Mutually exclusive pendulum preferences. We show the observed, and latent representations of the pendulum bob. (Top-bottom) The observed position and velocity of the pendulum bob; the learned representation of agent preferences on a 2-dimensional latent space, and the learned representation of agent preferences on a 1-dimensional latent space. In the 2-dimensional space, preferences are perfectly out of phase indicating that they are mutually exclusive.
Pendulum.
We model the pendulum as a single agent system represented as a graph with a single node. We construct a graphical representation at every timestep and set the node feature to the concatenation of the position and velocity of the pendulum bob.
The relationship between observed physical states and learned preferences is shown in Figure 4. Our model suggest the pendulum is a 1-dimensional system which is consistent with physical understanding. We can see this in two ways. First, since preferences $z_{1}$ and $z_{2}$ switch signs when the pendulum bob changes direction, we can interpret the learned categories as a desire to move clockwise ($z_{1}>0$) and a desire to move counter-clockwise ($z_{2}>0$). Moreover, the two preferences are perfectly out-of-phase and therefore mutually exclusive. This is observation is supported by consideration of the learned belief matrix whose entries satisfy the condition for mutual exclusivity, i.e., $a^{o}_{12},a^{o}_{21}\leq 0$,
$\mathbf{A}^{o}=\begin{bmatrix}0&-1.2947\\ -0.9147&0\end{bmatrix}.$ (19)
The conditions for mutual exclusivity are described in Section 3.1.
We empirically verify that the dynamics of this system can be captured with a 1-dimensional latent space. The prediction error for models with a 2-dimensional and 1-dimensional latent space are comparable i.e., $1.9718$e33-3- 3 MSE and $1.9916$e33-3- 3 MSE respectively. Qualitative results are shown in Figure 4 (bottom).
@@
In addition to providing a mechanism for identifying opportunities for dimensionality reduction, the nonlinear opinion dynamics inductive bias also provides a mechanism for controlling preference formation through the environmental input $b_{ij}$. We show trajectories with different rotational directions in Figure 5(a), and their corresponding environmental input latent preference pairs, ($b_{ij}$, $z_{ij}$), on the bifurcation diagram in Figure 5(b). The initial preferences for both trajectories are near neutral, but the initial environmental inputs are far from zero. The negative environmental input drives the initial preference of the blue trajectory toward an equilibrium in the preference space that corresponds to a strongly negative preference, and a clockwise motion in the physical space. The positive environmental input drives the initial preference of the orange trajectory toward an equilibrium in the preference space that corresponds to a strongly positive preference, and a counterclockwise motion in the physical space.
Double pendulum.
We use the double pendulum dataset to demonstrate interpretability in a more complex setting. We model the double pendulum as a two agent system represented as a graph with a two nodes. We construct a graphical representation at every timestep and set node features to the concatenation of the position and velocity of the corresponding double pendulum bob.
@@
Figure 6: Non-mutually exclusive double pendulum preferences. We show the observed, and latent representations of each double pendulum bob. (Top-bottom) The observed position and velocity of the first bob; the learned preferences of the first bob; the observed position and velocity of the second bob; the learned preferences of the second bob. When preference switch, physical bob’s motion switches direction.
@@
The relationship between the observed physical states and the learned preferences is shown in Figure 6. For the first bob, when the preference $z_{1,1}$ switches from greater (less) than to less (greater) than preference $z_{1,2}$, the physical bob switches from clockwise (counterclockwise) to counterclockwise (clockwise) motion. The opposite is true for the second bob. This behavior is observed across examples (see Appendix C.2). In contrast to pendulum preferences, double pendulum preferences are non-mutually exclusive. This observation is supported by the belief matrix, which does not satisfy the mutually-exclusive condition described in Section 3.1,
$\mathbf{A}^{o}=\begin{bmatrix}0&-2.8137\\ 1.1597&0\end{bmatrix}.$ (20)
@@ 5.2 Trajectory prediction
We demonstrate the efficacy of our model for trajectory prediction on several datasets. We report the quantitative performance of our model against competitive baselines in Table 1, and show the qualitative performance of our model in Figure 7. Dataset details are provided in Appendix A.
Table 1: Trajectory prediction. Trajectory prediction error on the Mass-Spring, Kuramoto, and TrajNet++ datasets is reported. Our BINN and BINN+ models outperform baseline models on all datasets.
Network Mass-Spring Kuramoto TrajNet++
BINN* $2.88\times 10^{-4}$ $4.52\times 10^{-3}$ $2.16\times 10^{-2}$
BINN+* $\boldsymbol{2.88\times 10^{-4}}$ $\boldsymbol{4.52\times 10^{-3}}$ $\boldsymbol{1.20\times 10^{-2}}$
NRI [19] $3.73\times 10^{-3}$ $8.20\times 10^{-3}$ $8.49\times 10^{-2}$
dNRI [20] $4.86\times 10^{-3}$ $6.41\times 10^{-3}$ $4.00\times 10^{-2}$
@@ 5.2.1 Mechanical systems
Mass-spring.
We demonstrate the utility of our model for trajectory prediction on the mass-spring dataset. The system consists of five masses and the dynamics are governed by a second-order linear equation. To model this system, we construct a fully-connected graph with five nodes and initialize node features as the concatenation of the position and velocity of the individual masses.
For both the BINN and BINN+ models we compute the communication matrix using the inter-agent distance defined in Equation (10). Our BINN models outperform both the NRI [19] and dNRI [20] models by an order of magnitude.
Kuramoto oscillator.
The Kuramoto oscillator dataset describes a 5-agent system with dynamics governed by a first-order nonlinear equation. To model this system, we construct a fully-connected graph with five nodes and initialize node features with the vector formed by concatenating the position and velocity of each oscillator.
For both the BINN and BINN+ models we compute the communication matrix using the inter-agent distance defined in Equation (10). On this dataset, our BINN models perform comparably with baseline models.
5.2.2 Pedestrian behavior
TrajNet++ simulator.
We evaluate our model on a synthetic subset of the TrajNet++ [55, 56] dataset where each trajectory contains 5 agents.
For our BINN model, we compute the communication matrix using the inter-agent distance as defined in Equation (10). For our BINN+ model, however, we assume pedestrian sensing scales inversely with inter-agent distance and introduce a prior $\tilde{\mathbf{A}}^{a}_{(i,j),t}$ on the communication matrix,
$\tilde{\mathbf{A}}^{a}_{(i,j),t}=\frac{1}{\lVert\mathbf{p}_{i,t}-\mathbf{p}_{i,t}\rVert^{2}+\epsilon},$ (21)
where $\epsilon$ is very small. We augment $\tilde{\mathbf{A}}^{a}_{(i,j),t}$ with a learned multiplier $\mathbf{A}_{(i,j)}^{\mathrm{pre}}$ which together determine the communication matrix,
Figure 8: TrajNet++ preferences. (Left) We visualize example trajectories from the dataset. The initial position of the trajectory is marked by a blue square and the point of direction change is marked by a black square. (Right) We plot agent preferences for each latent categories. Dominant $z_{i1}$ corresponds to leftwards motion and dominant $z_{i2}$ corresponds to rightward motion. Dominant $z_{i3}$ corresponds to downwards motion and dominant $z_{i4}$ corresponds to upwards motion.
@@ 5.2.2 Pedestrian behavior
On this dataset, our BINN and BINN+ models outperform competitive baselines.
The relationship between the observed trajectories and learned preferences is shown in Figure 8. Three consistent trends emerge across examples. First, when the relative magnitudes of $z_{i1}$ and $z_{i1}$ change, the direction of travel changes from right to left. Second, when the relative magnitude of $z_{i3}$ to $z_{i4}$ increases, agents increase their y-velocity. Finally, the dominant $z_{i3}$ in row 1 and 2 results in downwards motion, while the dominant $z_{i4}$ in row 3 results in upwards motion.
@@ 6 Conclusion
In this paper we address the challenge of designing a relational inference model with inter-agent relationships that are informed by preferences for multiple categories. To do this, we propose the use of nonlinear opinion dynamics as an inductive bias, and introduce an interpretable and flexible relational inference model. In addition to outperforming competitive existing relational inference models on multiple trajectory prediction datasets, our model provides a mechanisms for identification of mutually exclusive categories, and controlling preference formation.
Limitations and future work.
Similarly to prior work, our BINN models assume the same agents are present throughout a trajectory. Efforts to address this limitation would broaden the applicability of this work.

### v3

@@ 4.2 Latent nonlinear opinion dynamics
We use the nonlinear opinion dynamics formulation in Equation (1) to model the evolution of agent preferences on a set of latent categories,
@@
We learn the intrinsic agent parameters, $\mathbf{d}$, $\mathbf{u}$, and $\boldsymbol{\alpha}$, the extrinsic belief matrix $\mathbf{A}^{\mathrm{o}}$, and compute the inter-agent communication matrix $\mathbf{A}^{\mathrm{a}}_{t}$ at timestep $t$.
@@
where the timestep $\Delta t$ is dataset dependent.
@@ 4.3 Communication matrix
We define the communication matrix $\mathbf{A}^{\mathrm{a}}_{t}$ as a function of inter-agent proximity. In our baseline BINN model, we define $\mathbf{A}^{\mathrm{a}}_{t}$ as the distance between observed positions,
$a^{\mathrm{a}}_{(i,j),t}=\lVert\mathbf{p}_{i,t}-\mathbf{p}_{j,t}\rVert^{2},$ (11)
We also introduce Behavior-Inspired Neural Network+ (BINN+), where we incorporate an informed hypothesis of how agents in the system communicate. For example, for systems in which agents have more influence on each other when they are closer (e.g., human interaction) we define $\mathbf{A}^{\mathrm{a}}_{t}$ as the inverse distance between observed positions,
$a^{\mathrm{a}}_{(i,j),t}=\frac{1}{\lVert\mathbf{p}_{i,t}-\mathbf{p}_{j,t}\rVert^{2}+\epsilon},$ (12)
where $\epsilon$ is a small constant.
@@ 4.4 Decoder
Our decoding network $D_{x}$, is an MPNN that maps agent preferences to predictions of agent states. At every timestep $t$, the latent preferences of the multi-agent system are represented as a fully-connected graph with node values determined by the preferences of each agent.
Our decoder $D_{x}$ performs the following message passing functions for agent $i$ at timestep $t$:
@@ 4.5 Loss function
$\mathcal{L}_{\mathrm{latent}}=\frac{1}{T\mathcal{N}_{a}}\sum_{t=1}^{T}\sum_{i=1}^{\mathcal{N}_{a}}\Big{\lVert}\Delta\mathbf{z}_{i,t+1}-f_{\mathrm{NOD}}\left(\mathbf{z}_{i,t}\right)\Big{\rVert}^{2},$ (19)
where $\Delta\mathbf{z}_{i,t+1}=\left(\mathbf{z}_{i,t+1}-\mathbf{z}_{i,t}\right)/\Delta t$. This loss encourages alignment between the representations encoded by $E_{z}$ and those predicted by $f_{\mathrm{NOD}}$.
@@ 5.1 Interpretability of the latent space
Figure 4: Mutually exclusive pendulum preferences. We show the observed, and latent representations of the pendulum bob. (Top-bottom) The observed position and velocity of the pendulum bob; the learned representation of agent preferences on a 2-dimensional latent space, and the learned representation of agent preferences on a 1-dimensional latent space. In the 2-dimensional space, preferences are out of phase indicating that they are mutually exclusive.
Pendulum. We model the pendulum as a single agent system represented as a graph with a single node. We construct a graphical representation at every timestep and set the node feature to the concatenation of the position and velocity of the pendulum bob.
@@
The relationship between observed physical states and learned preferences is shown in Figure 4. Our model suggests the pendulum is a 1-dimensional system which is consistent with physical understanding. We can see this in two ways. First, since the preferences $z_{1}$ and $z_{2}$ switch signs when the pendulum bob changes direction, we can interpret the learned categories as a desire to move clockwise ($z_{1}>0$) and a desire to move counter-clockwise ($z_{2}>0$). Secondly, since the two preferences are perfectly out-of-phase and the entries of the learned belief matrix are negative,
$\mathbf{A}^{\mathrm{o}}=\begin{bmatrix}0&-1.2947\\ -0.9147&0\end{bmatrix},$ (20)
the conditions for mutual exclusivity is satisfied (see Section 3.1).
We empirically validate that the dynamics of this system can be captured with a 1-dimensional latent space. The prediction error for models with a 2-dimensional and 1-dimensional latent space are comparable i.e., $1.9718$e33-3- 3 MSE and $1.9916$e33-3- 3 MSE respectively. Qualitative results are shown in Figure 4 (bottom).
@@
In addition to providing a mechanism for identifying opportunities for dimensionality reduction, the nonlinear opinion dynamics inductive bias also provides a mechanism for controlling preference formation through the environmental input $b_{ij}$. We show trajectories with different rotational directions in Figure 5(a), and their corresponding environmental input latent preference pairs, ($b_{ij}$, $z_{ij}$), on the bifurcation diagram in Figure 5(b). The initial preferences for both trajectories are near neutral, but the initial environmental inputs are far from zero. The negative environmental input drives the initial preference of the blue trajectory toward an equilibrium in the preference space that corresponds to a strongly negative preference, and a clockwise motion of the pendulum bob in the physical space. The positive environmental input drives the initial preference of the orange trajectory toward an equilibrium in the preference space that corresponds to a strongly positive preference, and a counterclockwise motion of the pendulum bob in the physical space.
Double pendulum. We use the double pendulum dataset to demonstrate interpretability in a more complex setting.
@@
Figure 6: Non-mutually exclusive double pendulum preferences. We show the observed, and latent representations of each double pendulum bob. (Top-bottom) The observed position and velocity of the first bob; the learned preferences of the first bob; the observed position and velocity of the second bob; the learned preferences of the second bob. When the dominance of the preference switch, the physical bob’s motion switches direction.
We model the double pendulum as a two agent system represented as a graph with a two nodes. We construct a graphical representation at every timestep and set node features to the concatenation of the position and velocity of the corresponding double pendulum bob.
@@
The relationship between the observed physical states and the learned preferences is shown in Figure 6. For the first bob, when the preference $z_{1,1}$ switches from greater (less) than to less (greater) than preference $z_{1,2}$, the physical bob switches from clockwise (counterclockwise) to counterclockwise (clockwise) motion. The opposite is true for the second bob. This behavior is observed across examples (see Appendix C.1).
Table 1: Trajectory prediction. Trajectory prediction error on the Mass-Spring, Kuramoto, TrajNet++, and NBA Player datasets is reported. Our BINN and BINN+ models outperform baseline models on all datasets.
Network Mass-Spring Kuramoto TrajNet++ NBA Player
BINN* $\boldsymbol{2.88\times 10^{-4}}$ $\boldsymbol{4.98\times 10^{-3}}$ $2.16\times 10^{-2}$ $5.12\times 10^{-3}$
BINN+* $\boldsymbol{2.88\times 10^{-4}}$ $\boldsymbol{4.98\times 10^{-3}}$ $\boldsymbol{1.20\times 10^{-2}}$ $\boldsymbol{4.19\times 10^{-3}}$
NRI $3.73\times 10^{-3}$ $8.20\times 10^{-3}$ $8.49\times 10^{-2}$
dNRI $4.86\times 10^{-3}$ $6.41\times 10^{-3}$ $4.00\times 10^{-2}$ $6.42\times 10^{-3}$
In contrast to pendulum preferences, double pendulum preferences are non-mutually exclusive. This observation is supported by the belief matrix, which does not satisfy the mutually-exclusive condition described in Section 3.1,
$\mathbf{A}^{\mathrm{o}}=\begin{bmatrix}0&-2.8137\\ 1.1597&0\end{bmatrix}.$ (21)
@@ 5.2 Trajectory prediction
We demonstrate the efficacy of our model for trajectory prediction on several datasets. For each system, we construct a fully-connected graph with $\mathcal{N}_{a}$ nodes, and initialize node features to the concatenation of the position and velocity of the corresponding agent. We report the quantitative performance of our model against competitive baselines in Table 1, and show the qualitative performance of our model in Figure 7. Dataset details are provided in Appendix A.
@@ 5.2.1 Mechanical systems
Mass-spring. We demonstrate the utility of our model for trajectory prediction on the mass-spring dataset. The system consists of five masses and the dynamics are governed by a second-order linear equation.
We compute the communication matrices for our BINN and BINN+ models using the inter-agent distance defined in Equation (11). The predictive performance of our BINN models exceeds the performance of the NRI (Kipf et al., 2018) and dNRI (Graber & Schwing, 2020) models by an order of magnitude.
Kuramoto oscillator. The Kuramoto oscillator (Kuramoto, 1984) dataset describes a 5-agent system with dynamics governed by a first-order nonlinear equation.
We compute the communication matrices for our BINN and BINN+ models using the inter-agent distance defined in Equation (11). On this dataset, our BINN models perform comparably with baseline models.
5.2.2 Human behavior
TrajNet++ dataset. The TrajNet++ trajectory forecasting dataset (Kothari et al., 2021) is a collection of interaction-centric multi-agent datasets. We evaluate our model on a synthetic subset of the TrajNet++ dataset. This subset is comprised of 43,697 agent-agent centric trajectories. Each trajectory is $19$-timesteps long and consists of five agents. Since $\Delta t$ is not provided for the synthetic dataset, but is required by our model, we arbitrarily set $\Delta t=1$ for all experiments.
We compute the communication matrix for our BINN model using the inter-agent distance defined in Equation (11). For our BINN+ model, we assume pedestrian sensing scales inversely with inter-agent distance and introduce a prior $\tilde{\mathbf{A}}^{a}_{(i,j),t}$ on the communication matrix,
$\tilde{\mathbf{A}}^{a}_{(i,j),t}=\frac{1}{\lVert\mathbf{p}_{i,t}-\mathbf{p}_{j,t}\rVert^{2}+\epsilon}.$ (22)
We augment $\tilde{\mathbf{A}}^{a}_{(i,j),t}$ with a learned multiplier $\mathbf{A}_{(i,j)}^{\mathrm{pre}}$, and form the communication matrix by way of their product,
@@ 5.2.2 Human behavior
On this dataset, our BINN and BINN+ models outperform all baseline models. The relationship between the observed trajectories and learned preferences is shown in Figure 8. Three consistent trends emerge across examples. First, when the relative magnitudes of $z_{i1}$ and $z_{i2}$ change, the direction of travel changes from right to left. Second, when the relative magnitude of $z_{i3}$ to $z_{i4}$ increases, agents increase their y-velocity. Finally, the dominant $z_{i3}$ in row 1 and 2 results in downwards motion, while the dominant $z_{i4}$ in row 3 results in upwards motion.
Figure 8: TrajNet++ preferences. (Left) We visualize example trajectories from the dataset for a single agent. The initial position of the trajectory is marked by a blue square and the point of direction change is marked by a black square. (Right) We plot the agent preferences for each latent categories. Dominant $z_{i1}$ corresponds to leftwards motion and dominant $z_{i2}$ corresponds to rightward motion. Dominant $z_{i3}$ corresponds to downwards motion and dominant $z_{i4}$ corresponds to upwards motion.
NBA player dataset. The SportUV dataset (Linou et al., 2016) is comprised of 358,217 real-world NBA player trajectories. Each trajectory is 15 timestep long and consists of 10 agents.
We compute the communication matrix for our BINN model using the inter-agent distance defined in Equations (11), and the communication matrix for our BINN+ model using Equation (22) and (23). Our BINN$+$ model performs comparably to the dNRI model. The relationship between the observed trajectories and learned preferences is shown in Appendix C.2, Figure 10.
5.3 Robustness to choice of latent dimension
We investigate the robustness of our BINN model to the choice of latent dimension. We report the performance of our BINN model for large and small latent dimension in Table 2 . Across experiments, the latent dimension of BINN-large is twice that of BINN-small. The performance variation across models is minimal for all experiments, demonstrating the robustness of our model to the choice of latent dimension. We attribute this robustness to the nonlinear opinion dynamics inductive bias which imparts a mechanism for identifying mutually exclusive categories. We report the latent dimension used in each experiment in Appendix B, Table 5.
Table 2: Robustness to choice of latent dimension. We compare the trajectory prediction performance of our BINN model with different latent dimensions. We use a lower latent dimension in BINN-small and a higher latent dimension in BINN-large. The predictive performance across models is comparable.
Dimension Mass-Spring Kuramoto TrajNet++
BINN-small $\boldsymbol{2.88\times 10^{-4}}$ $4.98\times 10^{-3}$ $\boldsymbol{1.20\times 10^{-2}}$
BINN-large $3.33\times 10^{-4}$ $\boldsymbol{4.52\times 10^{-3}}$ $2.21\times 10^{-2}$
@@ 6 Conclusion
In this paper, we address the challenge of designing a relational inference model with inter-agent relationships that are informed by preferences for multiple categories. To do this, we propose the use of nonlinear opinion dynamics as an inductive bias, and introduce an interpretable and flexible relational inference model. In addition to outperforming competitive existing relational inference models on multiple trajectory prediction datasets, our model provides a mechanisms for identification of mutually exclusive categories, and controlling preference formation.
Limitations and future work. Similarly to prior work, our BINN models assume the same agents are present throughout a trajectory. Efforts to address this limitation would broaden the applicability of this work.
Acknowledgements
This work was partially supported by NSF grant MRSEC DMR-2011750 and the Princeton Catalysis Initiative. We sincerely thank the AISTATS 2025 reviewers and area chair for their thoughtful feedback.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model: Yes. Theoretical background for opinion dynamics is described in Section 3 and Appendix D and description of the algorithm is given in Section 4 and Figure 3.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm: Yes. Memory requirement for the proposed method and comparison with baselines are given in Table 4.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. Code is open sourced at https://github.com/CAB-Lab-Princeton/Behavior-Inspired-Neural-Network.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results: Yes. Assumptions we make about the nonlinear opinion dynamics model are given in Section 3 and Appendix 11.
    2. (b)
      Complete proofs of all theoretical results. Not applicable. Theoretical results are not part of the contribution to this paper.
    3. (c)
      Clear explanations of any assumptions: N/A.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL): Yes. All data generation and training hyper-parameters are given in Table 3 and 5. Code is open sourced.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen): Yes. Training details are given in Table 5.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times): N/A.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider): Yes. Information is provided in Appendix B. Specific information on the Neuronic cluster is availible at https://clusters.cs.princeton.edu.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets: Yes. TrajNet++ (Kothari et al., 2021) and SportsVU (Linou et al., 2016) are cited.
    2. (b)
      The license information of the assets, if applicable: N/A.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable: N/A.
    4. (d)
      Information about consent from data providers/curators: N/A.
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content: N/A.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots: N/A.
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable: N/A.
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation: N/A.